OpenStudy (anonymous):

Find the volume of a sphere that has a surface area of 16 sq. in.

hero (hero):

\[V_{sphere} = \frac{4\pi r^3}{3}\]

hero (hero):

\[SA_{sphere} = 4 \pi r^2\]

OpenStudy (anonymous):

(32/3)cubic inches 8cubic inches (4/3)cubic inches

OpenStudy (anonymous):

possible answers

hero (hero):

Substitute what is given into the surface area of a sphere formula

OpenStudy (anonymous):

ok

hero (hero):

\[16 = 4 \pi r^3\]

hero (hero):

Then solve for r

hero (hero):

Once you find r, substitute that into the Volume of a Sphere formula